### v1

# EcoServe: Designing Carbon-Aware AI Inference Systems

Abstract
The rapid increase in LLM ubiquity and scale levies unprecedented demands on computing infrastructure. These demands not only incur large compute and memory resources, but also significant energy, yielding large operational and embodied carbon emissions. In this work, we present two main observations. First, while GPUs dominate operational carbon, host processing systems (e.g., CPUs, memory, storage) dominate embodied carbon. Second, based on traces from production deployment of two Generative AI services in the cloud, offline, batch-inference accounts for a significant portion (up to 50%) of serving capacity. We propose four pillars of carbon-conscious infrastructure design for LLM serving systems: Reduce, Reuse, Rightsize, and Recycle. We demonstrate that EcoServe can lower carbon emissions by up to 47%, compared to performance, energy, and cost-optimized design points, while maintaining performance targets and SLOs.
I Introduction
The rapid growth of Large Language Models (LLMs) has led to unprecedented demands on computing infrastructure. From 2019 to 2022, we have seen an over 100$\times$ increase in model size from GPT-2 (1.5 billion parameters) to GPT-3 (175 billion parameters) to PaLM (540 billion parameters) [12]. The increasing compute demands have also led to dramatic increases in the power and energy demands for deploying LLM’s at-scale [67, 44]. Given the increasing demand and corresponding energy consumption, the Electric Power Research Institute (EPRI) projects that AI-driven data centers will account for up to 40% of the power consumed in local grids (i.e. Virgina, Oregon) [2]. The rising compute demand of deploying LLMs not only impacts the efficiency of the data center, but also the overall environmental impact of GenerativeAI and data centers. The increasing power and energy demands for LLM inference have led to significant research from the machine learning, systems, and computer architecture community. Systems such as vLLM [68], SGLang [75], Sarathi-serve [6], DynamoLLM [62], Melange [38], and Splitwise [53] have made significant strides in optimizing GPU utilization, throughput, cost, and energy efficiency via various intelligent resource provisioning, allocation, and scheduling strategies.
Figure 1: (Left) Breakdown of power (TDP) and embodied carbon between host systems (CPU) and GPU. (Right) EcoServe optimizes operational and embodied carbon based on input hardware, model, and workload traces by reusing CPU for offline inference, rightsizing GPUs, reducing underutilized hardware, and recycling hardware.
Although current approaches to LLM serving optimization focus on performance, power, energy efficiency, and cost, they do not directly address the growing environmental impact of LLM deployment. Crucially, recent work shows that the carbon footprint of AI is determined not only by operational emissions but also by embodied emissions [31, 71, 70]. While operational emissions are a product of energy consumption and carbon intensity of local power grids, embodied emissions owe to carbon impact from hardware manufacturing (i.e., application processors, memory, storage) data center construction. Recent academic and industry efforts to understand the environmental impact of data centers, as well as publicly available industrial sustainability reports show that more than 50% carbon emissions from hyper-scalars driven by renewable energy are due to embodied carbon [31, 39, 47, 27, 45]. Although significant attention has been paid to operational and energy efficiency, the embodied carbon footprint of hardware components remains largely overlooked.
Embodied carbon poses new trade-offs in the design space. Figure 1 illustrates the power and embodied carbon breakdown of two AI inference systems comprising NVIDIA A100 and H100 GPUs. While the GPUs account for around 75% of the thermal dissipated power (TDP), the host-processing system accounts for around 75% of embodied carbon. The high embodied overheads from host systems owe to a combination of memory capacity, storage, and mother-board (e.g., PCB, peripheral interconnects). While GPUs dominate the power consumption of AI servers, the embodied carbon is dominated by the host systems. Furthermore, during LLM inference, despite the CPUs being responsible for key operations like tokenization, graph computation, tensor management, and scheduling, they exhibit poor utilization; we observe around 6% average utilization per core based on characterization with vLLM (v0.6.2) on a 16-core server, effectively wasting available resources. Finally, embodied carbon of newer AI inference systems grows as they implement more compute, memory bandwidth, memory capacity, and storage.
In order to tackle these challenges, in this paper we present EcoServe, a carbon-aware resource allocation and scheduling framework for LLM serving. Central to EcoServe is co-designing decisions across capacity planning to minimize embodied carbon, resource allocation based on workload and heterogeneous hardware, and runtime scheduling to minimize operational emissions. We find separately considering these phases leads to sub-optimal outcomes i.e., improving energy efficiency can come at the expense of reducing overall carbon emissions. This misalignment is particularly pronounced for LLM serving workloads, where the compute-bound prefill phase and memory-bound decode phase demand different resources. To guide carbon-aware AI serving system design, we first develop an embodied carbon model, built upon existing models for individual integrated circuits [30, 24, 34], tailored for AI systems by considering the full host system (e.g., memory, storage, PCB, and peripheral circuits) and GPUs (e.g., memory, GPU die, peripheral ICs). We identify Host CPU as an untapped opportunity, and argue that optimizing the carbon footprint requires leveraging the capacity of CPU inference.
Using the model, EcoServe’s cross-layer approach focuses on four key sustainability-driven design principles: a) reusing existing CPU and host memory systems to increase AI cluster capacity for offline inference which accounts for nearly 30% of inference capacity based on traces from a production cloud provider, b) right-sizing host and accelerator systems for various model size and phases, and input sizes to balance performance, efficiency, and carbon, with hardware-software co-design, c) reducing wasted embodied carbon by eliminating unused resources for leaner hardware at no performance overheads, and finally, d) recycling host systems across generations of AI servers while still replacing accelerators at the current rate, to optimize collectively for efficiency and carbon. We quantify and demonstrate these solutions across a wide array of models and deployment scenarios with 1) open-source traces like ShareGPT [13] and Azure Function Traces [53]; 2) traces from a cloud service provider for two production LLM applications.
This paper makes the following contributions:
  • We introduce EcoServe, a carbon-aware framework for LLM infrastructure design that optimizes both operational and embodied emissions. Fundamentally EcoServe comprises four design strategies—Reuse, Rightsize, Reduce, and Recycle—motivated by insights from fine-grained embodied carbon models for AI systems and detailed profiling and analysis of inference performance and efficiency across diverse hardware and serving environments (e.g., LLMs, SLO targets, demand load).
  • Using EcoServe we show the individual design strategies save around 29%, 25%, 34%, and 41% carbon for reuse, rightsize, reduce, and recycle respectively, on a collection of open-source LLMs on heterogeneous hardware platforms, while maintaining SLO compared to a performance-optimized designs.
  • Putting the EcoServe optimizations together using an ILP formulation to co-optimize performance, efficiency, and carbon, we demonstrate 1.4x-2.2 $\times$ carbon benefits with minimal performance degradation for both online and offline workloads across a variety of LLM models and workloads.
II Background
Component Embodied kgCO2e Source(s)
SoC’s Tech. & Area dependent ACT [30], iMec [24]
DDR4/LPDDR5 0.29 / GB TechInsights [35]
GDDR6 0.36 / GB TechInsights [35]
HBM2 0.28 / GB TechInsights [35]
HBM3e 0.24 / GB TechInsights [35]
SSD 0.110 / GB R740 LCA [10] SCARIF  [34]
PCB 0.048 / cm2 12 layers Dell R740 LCA [10]
Ethernet card 4.91 Dell R740 LCA [10]
HDD controller 5.136 Dell R740 LCA [10]
Cooling 7.877 per 100W Scaled with TDP[34]
PDN / PSU 3.27 per 100W Scaled with TDP, Schneider [59]
TABLE I: Embodied carbon modeling strategy for each hardware component in AI inference systems.
Carbon Emissions in Computing Infrastructure Computing infrastructure contributes to carbon emissions in two primary ways: operational carbon and embodied carbon. Operational carbon refers to emissions from energy consumption during system operation, including direct power consumption by computing components and indirect emissions from cooling systems.
Embodied carbon, often overlooked, encompasses emissions from manufacturing, transportation, and eventual disposal of hardware components. Prior works indicate that for typical data center hardware, embodied carbon can represent over 50% of total lifetime emissions, with memory and storage being particularly carbon-intensive in their manufacturing process [69, 64, 31].
Recent work like ACT, SCARIF, FOCAL, LLMCarbon [30, 34, 22, 21] has highlighted the importance of considering both types of emissions. However, existing modeling approaches often do not look at fine-grained GPU carbon vs Host carbon footprint and treat these as separate concerns rather than considering their complex interactions with workloads in real-world systems.
LLM Serving Systems
State-of-the-art LLM serving systems like vLLM, TensorRT-LLM, AlpaServe, SGLang, Sarathi-serve, DistServe, and DeepSpeed [55, 38, 6, 76, 53, 75, 4, 41] have made significant advances in optimizing resource utilization LLM inference.
LLM inference typically has two distinct phases and two types of service level objectives (SLO): compute-bound prompt computation phase and memory-bound decoding phase. Online inference prioritizes low-latency responses for interactive tasks, while offline inference focuses on high-throughput batch processing, usually having 24hr SLO, presenting different optimization requirements. Additionally, the large model sizes and substantial key-value (KV) cache storage make LLM workloads memory capacity-bounded, further complicating resource management.
To address these challenges, these inference systems leverage GPU hybrid parallelization strategies, combining data, tensor, and pipeline parallelism to optimize computation across GPUs. Advanced memory management techniques, such as activation checkpointing, memory offloading, and paged attention, help alleviate memory pressure and improve resource utilization. Scheduler optimizations, including dynamic batching and chunked prefill processing, and disaggregating prompting and decoding phases, ensure high throughput while meeting latency constraints, enabling these systems to efficiently handle diverse LLM inference workloads [53].
While these systems primarily focus on performance metrics such as throughput and latency, recent works like DynamoLLM [62] and Melange [28] have brought attention to energy and cost efficiency of LLM inference. However, optimizing for energy consumption or cost does not directly optimize carbon impact.
III Understanding AI Systems’ Carbon Footprint
Figure 2: Trends in bit density (left) and embodied carbon footprint (right) across varoius DRAM memory technologies for 3 different manufacturers.
Based on recent efforts quantifying the carbon footprint of computing systems [30, 34, 21], the aggregate carbon emissions of AI systems is the combination of operational and embodied carbon as shown in the equation below:
$CF_{Task}=(P_{host}+P_{gpu})\times t\times CI+CF^{emb}_{host}\frac{t}{LT}+CF^{emb}_{gpu}\frac{t}{LT}$
where $P$ represents the power consumption of the host and GPU, $t$ represents the workload run-time, $LT$ represents the hardware lifetime, and $CF^{emb}$ represents the host and GPU embodied carbon. In terms of operational emissions, we measure the power consumption of host-processing systems including CPUs, DRAM, SSD, using a combination of Intel RAPL [17]; the power consumption of GPUs is empirically measured using NVML [49]. The carbon intensity of the different data center power grids is based on geo-temporal variance of power grid carbon intensities [5, 69]. To estimate embodied carbon, we leverage a combination of openly available product environmental reports and life cycle analyses from industry [10] and recently published, open-source tools to quantify the emissions from hardware manufacturing including the Architectural Carbon Modeling Tool (ACT), SCARIF, and GreenSKU [30, 34, 69]. In this section we describe how we extend current methods to study the embodied carbon of AI inference platforms including host systems, motherboards, and GPUs.
III-A Modeling Embodied Carbon of AI Systems
The embodied carbon footprint (CF) owes to a combination of components for the host systems and GPUs, as shown below:
$CF^{emb}_{host,gpu}=\sum_{k}^{\text{DRAM, PWB, PDN, Cooling, Storage, SoC Die,Chasis}}CF_{k}$
For host systems, we must account for the main CPU die, DRAM memory, SSD and HDD storage, the main printed wiring board (PWB), power distribution network (PDN), and server chassis. Similarly, for the GPUs we must account for the main SoC, DRAM memory, PDN, cooling (e.g., heat sink), and PWB. Table I describes the salient parameters in the embodied carbon model for AI inference systems. For many components including PDN and cooling (e.g., heat sink) given the lack of specific technology information we rely on publicly available product environmental reports from Dell [10]. Below we detail the methodology for estimating the embodied carbon of the remaining salient components.
III-A1 Integrated circuits (Application Processors)
To compute the embodied carbon emissions of application processors (e.g., CPU die, GPU die) we rely on recently published architectural carbon modeling tools [30, 34, 24, 21]. For application processors we input the process technology and die area into ACT [30, 24].
III-A2 Memory and Storage
Recent efforts from industry demonstrate memory and storage dominate the embodied carbon of Microsoft’s data center hardware [69]. To quantify the memory and storage embodied carbon, we consider both memory technologies and capacities. For storage, we estimate SSD emissions based on the Dell R740 life cycle analysis [10], yielding an average estimate of 0.110 kg CO2e per GB capacity; this is a conservative estimate compared to estimates of 0.160 kg CO2e per GB from academic surveys [64]. For memory, we extend existing ACT and SCARIF models by scaling wafer-scale DRAM data for different technologies from TechInsights [35, 66, 3] by publicly available DRAM bit-densities per technology from memory vendors (e.g., SK Hynix, Micron, Samsung) as shown in Figure 2. The trends show newer technology nodes, given higher bit-densities, exhibit lower embodied carbon per GB capacity. For the purposes of this work we estimate the DDR4/LPDDR5, GDDR6, and HBM2, HBM3e embodied carbon footprint as 0.29, 0.36, 0.28, 0.24 kg CO2e per GB respectively.
III-A3 Mainboard Printed Wiring Board (PCB)
To estimate the embodied emissions of the mainboard printed wiring board (PWB) we use the Dell R740 LCA [10]. Given the PWB area of the Dell R740, 1925 square centimeters (176 kg CO2), is significantly higher than modern server-class host systems and GPU’s, we scale the PWB area for specific systems. Thus, we estimate the embodied carbon of PWB to be: Total PCB $\text{Total PCB CF}=0.048\times Area_{PCB}$.
Figure 3: Trends in embodied carbon, power, and cloud cost for different generations of GPUs. As GPU performance increases (left to right), power consumption, cost and embodied carbon exhibit distinct trends.  [8, 9, 1].
III-B Characterizing Total Carbon
Based on the carbon models and runtime energy measurement, we can study the carbon breakdown of AI inference platforms.
Observation 1: Along with power capacity (TDP), embodied carbon of GPUs continues to rise as systems increase in compute and memory bandwidth.
Figure 3 analyzes the embodied carbon estimates of various NVIDIA GPUs, breaking it down into key components such as memory, printed circuit board (PCB), power delivery network (PDN), cooling systems, and system-on-chip (SoC). Figure 3 illustrates these components using stacked bars to represent embodied carbon, with the thermal design power (TDP) plotted on the right axis. We observe a clear trend: with newer GPU generations, computational performance, memory bandwidth, and energy efficiency improve. However, the embodied carbon also rises with increasing computational performance, advanced thermal solutions required, and improved memory capacity. For instance, compared to an NVIDIA H100, and NVIDIA L4 incurs 3$\times$ lower embodied carbon.
Observation 2: Host systems levy high embodied carbon costs on end-to-end AI systems, primarily due to the mainboards, memory, and storage components.
Figure 4: Embodied carbon breakdown of full inference systems available in cloud offerings from Azure and LambdaLabs [9, 1], varying the number and type of GPU. Host-processing systems account for over half of the embodied carbon in AI systems, owing largely to memory, storage, and mainboard overheads.
Figure 4 shows the embodied carbon breakown of end-to-end inference systems including both GPUs and host systems. The analysis highlights that the motherboard, DRAM, and storage constitute a significant portion of the embodied carbon impact. However, for AI inference workloads, host systems are frequently underutilized compared to GPUs, leading to a disproportionate contribution of their embodied carbon footprint relative to their actual usage during operations. This observation motivates exploring strategies such as resource sharing on the host (e.g., model multiplexing), improving utilization through CPU-based inference, extending the lifetime of host system components, and reducing the memory subsystem’s footprint (see Section IV).
Observation 3: As low-carbon-intensity energy sources increasingly dominate hyperscalar data centers, the carbon footprint of AI infrastructure is increasingly determined from the embodied carbon of both GPUs and host systems.
Figure 5 compares the embodied carbon versus the operational carbon per second across datacenters with different power sources. The host and GPU embodied carbon remained fixed, and the operational carbon footprint varies with the carbon intensity of energy sources. We show that for regions with high carbon intensity (Europe and US-East), the GPU operational carbon dominates the total footprint. As we move to cleaner energy sources, the operational emissions drop, leaving the embodied footprint to dominate. Additionally, the GPU operational carbon typically accounts for the majority of the total operational carbon, while the host system’s embodied carbon footprint dominates the total embodied carbon. This suggests different optimization approaches for operational and embodied carbon respectively: reducing operational carbon requires improving GPU energy efficiency, such as by optimizing workload allocation, minimizing idle power consumption, integrating clean energy sources, and leveraging more energy-efficient GPUs. On the other hand, reducing embodied carbon requires increasing host system utilization through reusing underutilized resources, extending hardware lifetimes, and improving resource sharing to amortize the embodied carbon cost of host components (see Section IV).
Figure 5: Embodied and operational carbon breakdowns of an AI inference system comprising an NVIDIA A100 GPU. We assume a 4-year lifetime running Llama-13B. In data centers and power grids with high degrees of renewable energy (e.g., hyper-scalars), embodied carbon dominates the overall carbon footprint of AI systems.
IV EcoServe Design
Figure 6: Overall EcoServe system diagram. EcoServe takes as input heterogeneous hardware resources, model architectures, and workload traces and optimizes the operational and embodied carbon. EcoServe co-designs four strategies—reusing offline CPUs, rightsizing GPUs, reducing wasted hardware, and asymmetrically recycling hardware—via an ILP formulation to balance performance, energy efficiency, and carbon. The output of EcoServe can be used by external scheduling, and resource provisioning and allocation frameworks to guide system design decisions.
In this section we propose EcoServe, a system that optimizes the holistic carbon footprint of AI systems. Figure 6 shows the overall EcoServe design architecture. EcoServe takes as input LLM models, workload traces and service-level objectives (SLO’s), and hardware platform specifications to reduce carbon footprint while maintain inference performance. EcoServe targets both operational and embodied carbon through four design strategies: reuse, rightsizing, reduce, and recycling. Ecoserve reuses already existing, but unused CPU’s in AI inference systems to process offline/batch workloads, increasing effective throughout and lower resource demand, thus lowering embodied carbon. EcoServe then rightsizes which GPUs across a heterogeneous set (e.g., L4, A6000, A100, H100) to provision for online and offline inference. Based on underutilized host compute, memory, and storage resources and their high embodied overheads (see Section III-B), we reduce wasted resources to further lower carbon footprint. Finally, we propose asymmetrically extending hardware lifetimes and recycling host processing systems and GPUs to balance operational and embodied emissions of AI inference systems. The four design strategies represent a combination of software-only (reuse), resource provisioning and scheduling (right-sizing and recycling), and hardware design (reduce) techniques.
Figure 7: Roofline models of an Intel Sapphire Rapids CPU with 112 cores compared to an NVIDIA A100. We overlay different operators in the decode (solid and dash-dot line) and prompt computation phase (dashed line) for Llama-3-8b with the maximum batch-size that CPU/GPU can hold with context length 2048. At large batch sizes, the GPU is capacity bound ($\star$) and the roofline shows reusing CPUs ($\star$) for offline decode computation is practical.
IV-A Reusing CPUs for Offline Inference
Recall, Section III-B shows host processing systems incur high embodied carbon overheads in inference systems. To maximize the utility of these resources, in this section we propose leveraging host systems, which typically go underutilized, to process offline inference. During auto-regressive inference, CPUs are responsible for the following operations: (1) waiting for incoming requests and processing HTTP and GRPC operations, (2) input tokenization, (3) prefix matching and request ordering, (4) allocating tensors, memory, and meta data for batches, and (5) inference post-processing including de-tokenization, speculative decoding, and beam search processing. Profiling several models (i.e., opt125m, Metallama-3-8B, Mistral and Mixtral) and request rates (2-32 requests per second from ShareGPT [13]) on vLLM (v0.6.2) and SGLang [75], we find average CPU utilization is around 6% of a 16-core processor (e.g., on average only a single CPU core is used). Thus, CPU processing is not on the critical path of recent LLM inference schedulers [68].
Identifying opportunities for using CPUs. Given the underutilization, we must identify which operations can be offloaded from GPUs to CPUs to maximize the benefit. To guide this decision, Figure 7 shows the memory bandwidth and compute gap of an Intel Sapphire Rapids 112-core CPU with an A100 40G GPU. Given the gap between memory bandwidth is lower between the two systems, we find low-arithmetic intensity operations such as attention scoring and decoding phase operators more suitable for CPU inference. Furthermore, the gap between GPU and CPU performance is decreased when running large batch operations given the GPU’s lower capacity. Figure 7 shows the maximal batch-size of 16 and 512 for GPU and CPU running at context length 2048 in FP16.
Figure 8: CPU offline inference requires carefully balancing parallelism degree (PD) and arithmetic intensity of Linear operators (AI, FLOPS/Byte). The run-time schedule can co-design the tile and workload slice dimensions with architecture decisions to optimize for offline performance.
Optimizing CPU performance via better parallelization. To optimize CPU offline inference performance, we must carefully balance the degree of thread-level parallelism exploited per operation, as well as the operational arithmetic intensity, compute requirements, and memory bandwidth requirements. This optimal configuration depends on hardware (e.g., AMX and AVX efficiency, memory system), model, batch sizes, and sequence lengths of input queries. The optimal partitioning strategy involves carefully configuring the thread-level parallelism and tensor tiling dimensions to maximize performance. Figure 8 shows the impact of tile-size on arithmetic intensity (AI) and parallelism degree (PD) based on an analytical performance model assuming a hidden dimension size of 4096. In LLM’s QKV Projection, given an (M, K) and (K, N) GEMM, we focus on parallelizing and tiling across $N$ and $K$ Furthermore, attention operators also constitute a large part of latency, especially for long context workloads, where CPUs exhibit more benefit [11, 65, 16, 15, 32]. Parallelization along the keys/values sequence length dimension, in addition to the batch dimension, maximizes memory bandwidth utilization across all CPU sockets.
Following the roofline analysis, while prompt computation is not desirable on the CPU, we can move the KV cache in a layered pipelined fashion from GPUs to CPUs to perform decoding. Optimizing these parallelism dimensions we find can improve CPU performance by up to 3.67$\times$ compared to a Llama.cpp baseline [25]; across an array of model sizes and sequence lengths, we observe a 1.4$\times$ improvement in performance (Figure 14). As CPU inference optimization is a nascent field, we expect future efforts leveraging sparsity and other optimizations can further improve CPU offline performance.
Offline inference demand: Given their low utilization, we propose using available CPUs to perform offline inference. Central to exploiting CPUs for offline inference are two key observations: demand for offline versus online inference, and dinstinct SLOs for online and offline inference. Figure 9 illustrates the ratio of demand for offline and online inference for two services (A and B) in a real, production cloud provider over a week (left) and within a day (right). On average, offline inference accounts for 21% and 45% of total demand and infrastructure capacity for Service A and B, respectively. Note, at some points in the week, the offline demand peaks at around 27% and 55% for Service A and B respectively; shaving resource demands at this peak can reduce resource requirements. By leveraging available, underutilized CPUs and DRAM in existing systems, we reduce infrastructure capacity needed for offline inference given their relaxed SLOs.
Since offline workloads do not have stringent SLO, it is beneficial to run large batches on CPU and reduce the GPU provisioned. In practise, we can also offload offline demands back to GPU once the online GPU servers are idle for high carbon intensity regions, since GPUs are more energy-efficient. This runtime scheduler decision can be jointly designed with workload partitioner.
Figure 9: Breakdown of online and offline demand for two LLM services (A and B) running in a production, cloud data center over a week (left) and during a day (right). Offline demand account for a significant portion of overall Generative AI serving load, which can be processed by less performant hardware (e.g., CPUs).
Load-aware reuse: Figure 10 shows the server capacity (overprovisioned by 30%) needed for offline infrastructure to serve Llama-3-8B models under time varying demands based on production loads for Service A shown in Fig 9. We show the impact of using CPUs to either process offline inferences during peak demand (red) or continuously at all times (blue). We assume the provision reallocation happens every 4 hour. EcoServe can use CPUs from nodes where GPUs are processing both offline and online inferences to gain significant infrastructure capacity. Reusing CPUs during peak demand, we find can reduce offline capacity demands by nearly 1.32$\times$, saving significant embodied carbon at-scale. The capacity reduction shown is conservative, assuming fixed batch sizes for CPUs and GPUs (for shorter context length, larger models); maximizing batch size on CPUs can further reduce capacity by up to 45%.
IV-B Right-sizing GPU Provisioning
In this section we investigate provisioning heterogeneous GPU and scaling based on varying resource demands to optimize not just performance and efficiency, but also carbon footprint. Fundamentally, recent work has shown different model architectures, sequence lengths, and LLM phases (i.e., prompting and decoding) have different compute, memory bandwidth, and memory capacity requirements [62, 53, 52, 76]. Furthermore, modern GPUs have diverse hardware characteristics that impact performance, performance-per-watt, and performance-per-carbon tradeoffs.
Heterogeneous GPU partitioning. Figure 11 compares the energy efficiency (bottom) and embodied carbon efficiency (top) of NVIDIA A100 versus H100 GPU’s for Gemma 27b across different prompt lengths and batch-sizes. The analysis separately considers LLM prompt computation and decode (i.e., token generation) [53]. Values greater than one represent scenarios where the NVIDIA A100 GPU is preferred. In terms of prompt computation, we find the A100 is preferred for smaller input sequences and batch-sizes; as input sequence length and batch-size increase, H100 is more energy and carbon efficient. However, in terms of decode, the NVIDIA A100 is always better, owing to the low Model FLOPs Utilization (MFU) and Memory Bandwidth Utilization (MBU) of the hardware on H100 and high embodied carbon and energy cost of H100.
Figure 10: Applying EcoServe reuse optimizations, we find total infrastructure capacity for offline inference can be reduced by up to 45%. Red and blue curves show two strategies’ capacity provision lines by intelligently leveraging CPU cores to do offline inference.
To optimize energy and carbon efficiency of LLM serving, we propose a workload-aware co-design strategy for GPU provisioning that optimizes resource allocation based on workload characteristics. Contrary to traditional wisdom of provisioning based on the peak demand, we allocate GPU resources tailored to each workload slice’s characteristics in terms of performance and carbon efficiency, guided by offline profiling and performance modeling. Our modeling specifically distinguishes between prompt computation and decoding phases to better capture heterogeneity across hardware.
Exploiting tensor-model parallelism. The decision of how many levels of tensor parallelism (TP) is dependent on the target metrics. From latency perspective, sharding tensors across GPUs with high bandwidth communication links enables faster computation and more HBM bandwidth, it’s generally favorable for large models to do higher level of TP (up to the tile and wave quantization effect [50]). However, the latency comes at a cost with more communication overheads. From carbon perspective, the TP level depends on the ratio between CPU embodied and GPU embodied carbon. Table II details the desiderata for TP level deployment. In practice, we tailor it to each hardware, model, batch size and SLO.
Figure 11: Comparison of prompt computation (left) and decode (right) energy efficiency and carbon for Gemma-27b model on NVIDIA A100 and H100 GPUs. The results are normalized to A100 (higher than 1 means A100 is preferred). We find the optimal GPU varies based on LLM phase, prompt length, and batch-size, motivating the need for heterogeneous GPU rightsizing.
Metric Relative Scale Decision Criteria
Power $\frac{2nP_{GPU}+P_{CPU}}{nP_{GPU}+P_{CPU}}$ Consider if $P_{CPU}\gg nP_{GPU}$
Latency $\approx 0.5+C_{comm}$ Favorable when model $>$ memory
Cost/Model $\approx 1$ Near-constant when $C_{CPU}\ll nC_{GPU}$
Carbon $\frac{CF^{emb}_{CPU}+nCF^{emb}_{GPU}}{CF^{emb}_{CPU}/2+nCF^{emb}_{GPU}}$ Better with higher $CF_{CPU}/CF_{GPU}$
Energy $\approx 0.5$ Linear improvement with fixed CI
TABLE II: Power, latency, carbon, energy and cost when scaling up model and tensor-model parallelism.
[…]
IV-C Reduce: Host System Provision
This section examines strategies to minimize CPU resource underutilization to reduce embodied carbon waste. We focus on optimizing memory and storage resources, which comprise 36% of embodied emissions in Azure A100 offerings [^1] Standard_ND96asr_A100_v4, the number is generally higher for higher capacity instances. These optimizations must be balanced against the benefits of CPU reuse for offline inference, as explored in Section IV-A.
IV-C1 DRAM Reduce Strategy
DRAM supports AI inference through:
  • Memory offloading and intermediate data storage
  • Data processing and memory allocation
  • KV Cache swap space and prefix caching
For online inference with tight SLOs, offloading is not viable for large models. The minimum DRAM capacity is determined by:
$\begin{gathered}C^{\text{DRAM}}=C^{\text{Weight (layer)}}+C^{\text{KV/activation offload for online}}+C^{\text{KV for offline}}\\ \min C^{\text{DRAM}}=M_{kv}(n)=4ndh_{kv}l\end{gathered}$ (1)
where $d$ is model dimension, $l$ is layer count, $n$ is P90 aggregated context length with zero reuse distance, and $h_{kv}$ is KV head dimension. To mitigate performance impact from reduced DRAM, we use profiling to understand prefix reuse distances and determine optimal CPU cache space. We route longer-context requests to DRAM-abundant servers and leverage solutions like CacheGen [42] to offload caches across multiple instances or to SSD.
IV-C2 SSD Reduction Strategy
While cloud GPU offerings typically include large SSDs, this isn’t fundamental for inference serving. SSDs consume approximately 2.8 W per TB idle power (¿10% of server idle power). The minimum required SSD size approximates GPU memory capacity:
$\begin{gathered}C^{\text{SSD}}=1.2C^{\text{GPU}}+C^{\text{Model buffer}}+C^{\text{KV Offload for offline}}\\ \min C^{\text{SSD}}=1.2C^{\text{GPU}}\end{gathered}$ (2)
When SSD capacity is used to download and maintain models, we imagine remote storage with GPUDirect can reduce over provisioned storage.
IV-D Recycle: Asymmetric Lifetime Extension for Host and Accelerator
Due to the nature of fast energy efficiency improvements of GPU and slow energy improvement on hosts, the optimal lifetime $T$ is also different, necessitating an asymmetric recycling strategy.
IV-D1 Host Recycle
Extending life time or using older DRAM, SSD and CPU components can be helpful for reducing both the embodied and operational carbon. This opportunity of recycling arises from a key observation: the bandwidth improvements in memory and storage subsystems have not scaled proportionally with CPU/GPU performance across recent technology nodes. Our measurements across three generations of server hardware show that DRAM and SSD bandwidth often do not constitute the primary bottleneck in modern inference serving systems. This bandwidth headroom creates an opportunity to integrate older memory and storage components while maintaining target performance levels.
The key challenge is to make sure that the recycle strategy doesn’t impact performance SLO and the reliability of the serving system [60, 46, 61, 36]. From production systems, we find that the utilization of those components in AI inference systems is typically low, which means recycling and extending lifetime would not significantly impact reliability [69, 60].
IV-D2 GPU Recycle
The GPU recycle strategy depends on 1) inter-generational GPU energy efficiency improvement, 2) the workload and model characteristics, and 3) the upfront embodied carbon cost. For example, in Figure 12, although upgrading V100 to GH200 would pay an upfront embodied carbon cost, the new hardware would enable much more tokens to be processed per operational carbon. Hence the longer we use the newer GPU, the more relative carbon footprint saving we have, especially under a datacenter with less renewable energy source. Figure 12 demonstrates the optimal years of usage for different GPUs under different carbon intensity region for better carbon savings under a 70% maximum traffic load.
Figure 12: Relative carbon savings with various hardware compared with V100. (Left) prompt heavy workload with Carbon Intensity (CI) 400 gCO2e/KWh, (Middle) prompt heavy workload or (Right) decode heavy with CI = 27 gCO2e/KWh, The break-even points are different.
IV-E Putting it all together
Deploying EcoServe at scale requires careful consideration of multiple objectives, including cost, reliability, carbon, energy, and performance. Some of the strategies can conflict with each other. For example, for Reuse and Reduce, there is a balance one needs to strike, like how many numbers of CPU cores and DRAM/SSD minimum capacity value. These capacity planning phase decisions can be codesigned with runtime scheduling. As a hyperscale data center, it is common to have different backend servers with a variety of configurations, and we propose to use a gradient of different carbon optimization techniques (e.g., some Reuse heavy with larger memory subsystems, whereas others Reduce). A load balancer, informed by these capacity planning decisions, can dynamically perform runtime placement and KV cache transfers. This approach enables optimized trade-offs, such as improved performance-per-TCO or performance-per-carbon-efficiency, ensuring scalability and sustainability across diverse workloads.
Workload Slicing and Disaggregation: We represent the workload in two distinct phases: prompt and decode. Let $H(i,o)$ represent the request rate for inputs of length $i$ and outputs of length $o$. We put workload into histogram bucket $b$, and further divided into $S$ slices for fine-grained allocation, where slice factor $f$ determines the granularity. For a bucket with request rate $\lambda_{b}$, each slice $s$ has rate $\lambda_{s}=\lambda_{b}/f$.
For each slice $s\in S$, we define $\lambda_{s}$ as the request rate, $\text{Lat}(s,g,l,\phi_{s},m_{s})$ as expected latency for the prompt or decode phase on GPU $g$ with allocated resources at load $l$, required CPU cores $\phi_{s}$, and memory $m_{s}$. The load imposed by slice $s$ on GPU type $g$ is calculated separately for each phase:
$\footnotesize\text{Load}_{p/d}(s,g)=\frac{\lambda_{s}}{\text{MaxTput}_{p/d}(g,\text{size}(s),\text{SLO})}$
where $\text{size}(s)$ represents the input/output length pair of slice $s$.
Resource Model: For each GPU type $g\in G$, we define $c_{g},c_{\phi},c_{m}$ as hourly GPU, CPU, memory cost, $\gamma_{g}(t)$ as GPU carbon intensity at time $t$, which is both linear with operational power of the workload times the carbon intensity, and with unit-time embodied carbon of the hardware. The carbon impact of a workload is the sum of both phases $\text{Carbon}(s,g,l,\phi_{s},m_{s})=\sum_{k\in\{p,d\}}\gamma_{g}(t)\cdot\text{Lat}_{k}(s,g,l,\phi_{s},m_{s})$.
Decision variables include GPU assignment matrix $A\in\{0,1\}^{N\times M}$, GPU count vector $B\in\mathbb{Z}_{\geq 0}^{M}$; CPU cores and memory allocated to slice $s$ as $\Phi_{s},M_{s}$.
ILP Formulation: Our goal is to minimize the weighted carbon cost and latency-dependent hardware cost under service level objectives (SLO) based on both TTFT (time to first token) and TPOT (time per output token).
$\min_{A,B,\Phi,M}(1-\alpha)$ $\Big{[}\sum_{j=1}^{M}B_{j}c_{j}+\sum_{s=1}^{N}(\Phi_{s}c_{\phi}+M_{s}c_{m})\Big{]}+$
$\alpha\sum_{s=1}^{N}\sum_{j=1}^{M}A_{s,j}\text{Carbon}(s,j,l,\Phi_{s},M_{s})$
s.t. $\forall s,\quad\sum_{j=1}^{M}A_{s,j}=1$
$\forall j,\quad\sum_{s=1}^{N}A_{s,j}\cdot(\text{Load}_{p}(s,j)+\text{Load}_{d}(s,j))\leq B_{j}$
$\sum_{s=1}^{N}\Phi_{s}\leq\Phi,\quad\sum_{s=1}^{N}M_{s}\leq M\,\quad\forall s,\quad\Phi_{s}\geq\phi_{s},\quad M_{s}\geq m_{s}$
$\forall s,j\text{ where }A_{s,j}=1,\quad\text{Lat}_{p/d}(s,j,l,\Phi_{s},M_{s})\leq\text{SLO}_{p/d}$
where $\alpha\in(0,1)$ is a weighting variable between cost and carbon objectives. Unless otherwise stated, we choose $\alpha=1$. A smaller value such as $\alpha=0$ will be simply optimizing for cost. We build our carbon model $\gamma_{g}$ based on the runtime energy measurement from Intel RAPL or NVML libraries, GreenSKU for the carbon intensity [69, 49, 17], and our embodied carbon model.
V Implementation and Experimental Setup
Hardware Systems. The experimental evaluation uses a heterogeneous set of CPU and GPU hardware. In terms of GPUs we use a combination of PCIe-based NVIDIA H100, A100, A6000, L4, and A40 GPUs, representing a diverse range of compute, memory, power, and energy-efficiency tradeoffs. In terms of CPUs, we use a dual-socket server class Intel Sapphire Rapids CPU. The CPU frequency governor is set at performance mode with turbo boost disabled.
Models, Workloads and Load Generator. We evaluate our system using a wide range of workloads to capture diverse scenarios. For models, we assess different architectures with varying sizes. For request generator, we use scaled Azure function traces (AZF) and production offline workload traces [48] (Figure 9). For workload size, we sample from public datasets listed below:
[…]
Software. We use vLLM v0.6.2 and Llama.cpp (d5ed2b9) to deploy LLM instances across provisioned GPU and CPUs. We implement the profiling-based performance, energy and carbon modeling using Python based on those frameworks under various hardware, workload traces and request generators. We implement the ILP formulation and resource allocation algorithms based on CVXpy [19].
Carbon intensity (CI). We assume a CI of 27 gCO2e / kWh (aligned with US South in Azure) in the following data representation when we show the carbon metric, unless otherwise stated. EcoServe is not fundamentally tied to optimizing a single CI value and is able to optimize across different CI regions.
VI Evaluating EcoServe
We evaluate EcoServe across three dimensions: carbon efficiency, performance, and energy. Our evaluation focuses on the following key research questions:
  • How much additional operational and embodied carbon can EcoServe save in comparison to performance, energy, and cost-optimized baselines?
  • What are the carbon footprint implications of combining multiple categories of EcoServe’s optimization strategies? What are the tradeoffs with performance?
  • What is the effectiveness of individual design strategy in 4Rs?
VI-A End-to-end Evaluation
Figure 13: Carbon vs. performance trade-off. Lower-right is better. (Left, Center) We compare EcoServe to performance, energy, and cost (Melange) optimized baselines for prompt latency (TTFT, center) and generation latency (TPOT, left). We show the individual improvements (i.e., reuse, reduce, rightsize, recycle) as well as the aggregated benefit. EcoServe achieves up to 47% carbon reduction at comparable performance. (Right) Under different workload mixes, we show the cumulative benefits of successively applying EcoServe’s optimizations to online and offline capacity. All experiments assume iso-throughput.
We begin by summarizing the overall carbon and performance implications of EcoServe. Figure 13 (left) demonstrates the average relative performance degradation and carbon improvements normalized to a performance optimized (perf-opt) system configuration and resource allocation baselines. The baselines are configured as follows: (1) a performance optimized (perf-opt) configuration implements a single hardware that minimizes TTFT and TPOT, (2) a cost-optimized baseline (Melange) which provisions hardware to minimize cloud cost, (3) an energy-optimal configuration (energy-opt) that find configurations minimize energy. The workloads we evaluated on spans a variety of SLOs with the mixture percentage sampled from the production traces. We drive the workload with a Poisson request generator with various arrival rates and scaled Azure Function Traces (2023) to emulate bursty behavior of online samples (5% - maximum system load). The prompt and decode length distribution is sampled from ShareGPT [13].
Figure 13 (left,middle) shows that EcoServe variants can achieve significant carbon savings compared with performance optimal configurations for both online and offline workloads. We further decompose the carbon savings from various strategies:
  • Reduce can yield around 16-42% carbon savings. For leaner GPU offerings like T4, the savings are less than higher-end GPUs since the host is designed to scale with the GPU memory capacity.
  • Right sizing (RS) enables around 15-39% better carbon, leveraging the SLO slacks for various workloads on TPOT and TTFT. The benefit is particularly significant in lower carbon intensity regions with higher load variability and higher online workload component, as EcoServe can dynamically adjust its resource allocation to match demand to different workload slices.
  • Optimized CPU reuse strategies can save around 25-30% of carbon for clusters with different percentages of offline workload.
  • When it comes to combining three strategies together, RS + Reuse are software solutions that produce more synergy, whereas Reduce and Recycle are orthogonal hardware design optimizations. On average EcoServe yields 47% carbon savings.
Figure 14: Beginning with llama.cpp as a baseline, we optimize CPU attention for Gemma-2b and Gemma-27b GGUF models. EcoServe’s CPU decode optimization improve performance by up to 4.03$\times$ and on average by 1.34$\times$ (across batch-sizes and parallelism dimensions) on an Intel Sapphire Rapids 56 core CPU.
VI-A1 Performance Analysis
From Figure 13, we demonstrate that the performance of EcoServe variants is close to Perf-opt within 3 percent, except for the reuse case. The reason for a worse TPOT (around 5%) is because of the offline workloads, where the reuse strategy prioritize the decoding of tokens with longer context length to be offloaded to CPU. This is not perf-optimal when there are further GPU capacities to use for serving offline workloads.
EcoServe (RS) provides better improvements with TPOT metric than TTFT metric in the presence of SLO slack. This is because of the longer prompt and higher load, when we are resizing the allocation, the unavailability will cause the prompting scheduler to temporarily queue the unserved requests’ prompts, and the queuing delay increases TTFT. Furthermore, by design, our Reuse mostly addresses the decoding servers’ capacity planning with CPUs, hence the performance on TTFT on the prompting servers is untapped. EcoServe (Reduce) incurs a little overhead on TPOT (¡5%); but energy-opt allocator is worse in performance through aggressively choosing energy-opt hardware, ignoring the embodied carbon.
VI-A2 Breakdown Analysis
The benefits of carbon footprint reduction in EcoServe can be uncovered layer by layer through various strategies (Figure 13 Right).
For online workloads, Reduce (online resource) allows us to save approximately 17-40% of the carbon footprint by reducing DRAM and SSD capacity. When moving to Reuse, online workloads do not use CPU resources, maintaining consistent savings. The RS (online) strategy further increases savings by employing workload-aware load balancers and resource allocators, limiting the overprovisioning of high-end hardware.
For offline workloads, savings from the Reduce strategy are less significant. The Reduce (offline resource) method is conservative, requiring DRAM to match the model size for offline inference on CPU. However, substantial savings come from reusing offline CPU capacity for decoding tasks. Rightsizing also yields significant benefits due to relaxed SLO and the Integer Linear Programming (ILP) formulation to optimize GPU sets through disaggregation. Longer offline workloads yield more benefits.
For mixed workload production traces, workload (A), which has less offline demand, exhibits less benefit from aggressive Reduce (online); however, the overall carbon savings stay relatively similar for the two mixed workloads based on our analysis, demonstrating the generality for our approach across different types of workloads.
VI-B Reuse Strategy Evaluation
We evaluate the two design aspects of the reuse strategy: first, optimized CPU threading and tiling implementation for GEMM and long-context decoding. Second, workload- and load-aware CPU reuse strategy. Figure  14 shows the speedup with various workload sizes with EcoServe’s optimized reuse strategy, compared with state-of-the-art CPU inference engine baseline (llama.cpp) on Gemma 2/27B. The evaluation varies the core counts (blue as 56 and black as 112) to show the generality of the approach. Through choosing the right tiling and parallelism strategy guided by performance modeling, we improve decoding latency, which directly translate to higher throughput and carbon savings of 1.34x on average.
Figure 15: Decoding throughput (Left), operational (Middle) and embodied carbon (Right), for various models and workload sizes, comparison with naive llama.cpp vs. optimized CPU reuse strategy, normalized to A100 GPU performance under maximum throughput. Carbon reduction is 12, 26% for Gemma-27B, and 40%, 66% for Llama-8B respectively under short and longer workload.
Figure 15 analyzes the detailed embodied and operational carbon breakdowns compared with serving on GPUs for EcoServe’s reuse, showing the benfit of workload- and load-aware optimization. Given the CPU’s lack of energy proportionality, the added operational power is relatively smaller, but still results in operational carbon cost with longer execution time. The free lunch from a 56-core SPR machine CPU attached to the A100 enables us to have 0.53-2.29$\times$ of the throughput of the GPU (Blue bars). Although CPU generally has worse energy efficiency, especially for larger models when both CPU and GPU are compute bound (Gemma, Orange bars) EcoServe’s CPU decode can achieve up to 66% relative embodied carbon savings compared with a naive llama.cpp (CPU-Naive) strategy as 48%. On average, EcoServe can achieve 3.51x embodied carbon savings compared with non optimized baseline across the above workloads. More importantly, without EcoServe’s reuse strategy (CPU-Opt), the embodied carbon for llama.cpp (CPU-naive) is worse than non-reuse (GPU) for Gemma-27B model and Llama-8B model short context cases. This is due to the fact that under longer context and smaller model dimension, the gap between CPU and GPU throughput is larger due to batching effect, and the carbon saving is larger in an iso-throughput setting.
The total carbon benefit comes from the net effect of embodied and operational carbon with Reuse. Although the operational carbon is not better than GPU, by leveraging the unsused CPU cores, EcoServe improves both in cost and embodied carbon; fewer GPUs need to be over-provisioned. Furthermore, from throughput perspective, EcoServe demonstrates more benefits for long context decode with customized CPU kernel optimization, and we can enable flexible carbon reduction with runtime scheduling based on workload sizes and demands.
Figure 16: Relative carbon and energy improvements of rightsizing GPUs on Gemma-27B on homogeneous GPUs (i.e., H100, A100, L4), and cost-optimized provisioning (Melange). EcoServe’s Right-sizing strategy exploits LLM phases to reduce overall energy and carbon under TPOT = 100 ms (online) and TTFT = 24hr offline constraints.
VI-C Right-sizing Strategy Evaluation
For right-sizing, we evaluate our system on Gemma-27B with various load and workload against the following two types of baselines:
  • Melange: Cost-optimal GPU resource allocation.
  • Single hardware: pick the single hardware and its replica numbers to satisfy workload demands that disregard SLO slacks and the energy/carbon implication for hardware choices.
Figure 16 shows the final outcomes for right-sizing strategy given different workload types and load of the system. Each request generator runs for 5 minutes length traces, and we discard the first and last 5% of the statistics for warming up effects. The length of the workloads are randomly sampled from the open-sourced traces with both long and short prompts.
For online setting (left), under lower request rate, EcoServe consistently outperforms Melange by over 2.56x for carbon and 2.25x for energy. The improvement owes to separate the resource allocation decisions for LLM phases; finer-grained allocation enables better utilization and energy. As an example, for request rate equal to 1, EcoServe will choose the most carbon efficient hardware based on the workload size; e.g. for medium-sized workload it chooses L4 for decoding and A100 for prompting; for long prompt it chooses A100 for decoding and H100 for prompting. However, Melange sticks to the best perf-per-cost hardware for both phases for each workload slice, and the quantization effect for available hardware capacity introduces more over-provisioned resources for decode phase, which negatively impacts embodied and operational carbon.
EcoServe also performs better than other single hardware choices both on carbon and energy. The closest baseline is L4 due to its higher energy and carbon efficiency. This is because of the flexibility in allocating a combination of resource based on the optimal energy or carbon profile for different workload slices under various SLO.
Furthermore, for offline setting (offline), EcoServe consistently wins in both energy and carbon up to 1.6$\times$ with the next best baseline, as it considers both dimensions in the optimization objective. Under the carbon intensity of US-Central / South, operational carbon is around 0.64x of the embodied, making EcoServe a desirable solution for balancing both energy and embodied carbon.
VI-D Recycle Strategy Evaluation
Figure  17 shows the absolute carbon cost (top) and relative carbon savings (bottom) across years with a fixed (baseline) and asymmetric (EcoServe) upgrade strategy. The baseline server configuration assumes an embodied carbon of 800 kgCO2e for the host and 120 kgCO2e for the GPU. Additionally, the operational carbon emissions per year for the server are set at 600 kgCO2e. We assume that hardware GPU energy efficiency doubles every 3.5 years [63]. In the baseline fixed upgrade strategy, both the host systems and GPU accelerators have the same 4-year upgrade cycle. EcoServe, however, proposes an asymmetric upgrade strategy where host systems are upgraded less frequently (every 10 years) while GPUs are upgraded more often (every 3 years) to better align with their energy efficiency improvements.
The bottom plot demonstrates the relative carbon savings enabled by the asymmetric strategy. By upgrading host systems less frequently, significant embodied carbon savings are achieved due to their higher upfront embodied carbon costs. Although this can cause some negative operational carbon savings, host systems have less impact on overall operational emissions to AI inference compared to GPUs. At the same time, more frequent GPU upgrades, which incur relatively lower embodied carbon costs, allow for earlier adoption of energy-efficient hardware, leading to operational carbon savings. This trade-off enables EcoServe to balance embodied and operational carbon reductions. Over the 10-year period, EcoServe asymmetric strategy is able to achieve approximately 16% cumulative carbon saving compared to the baseline fixed strategy.
Figure 17: Trends in annual embodied and operational carbon based on EcoServe’s assymetric recycling of CPUs (10 years) and GPUs (3 years). Baseline is a fixed schedule for both CPU and GPU (4 years). Assuming energy efficiency doubles every 3.5 year [63].
VII Related Work
Existing CPU Inference LLM inference on CPUs, as demonstrated by frameworks like Llama.cpp and Gemma.cpp, has garnered significant attention for enabling cost-effective deployment [25, 26]. These systems employ optimizations such as quantization and memory-mapped weights to run models with limited resources. Recent studies also focus on improving CPU-aware inference efficiency through sparsity-aware techniques, quantization optimizations, offloading, and library enhancements (e.g., OpenBLAS [18], oneDNN [51], XNNPACK [56]), some of which are implemented in llama.cpp [25]. We built our work on top of llama.cpp and come up with new solutions for codesigning resource provisioning and workload scheduling.
Heterogeneous LLM on Commodity Hardware: Recently, people have gained more interest in deploying or training foundational models under decentralized or heterogeneous commodity setups [72, 58]. To account for the heterogeneity in the interconnect bandwidth of commodity setup, Mobius [23] introduces a novel pipeline parallelism scheme enabling heterogeneous memory for large-scale model training, while bringing fewer communications than existing systems. While those works mostly focus on model partition and placement, we focus more on the request scheduling and resource provision. Our objective which includes carbon is also fundamentally different from performance-oriented systems.
CPU Memory Offloading Recently, memory offloading has gained popularity as a research area due to the limitations and high costs of GPU HBM. DeepSpeed ZeRO-Inference [7] can run LLMs with offloading. Specifically for MoE models, Huang et al [33] propose to swap experts to CPU memory to reduce the GPU memory consumption. PC-MoE [37] proposes to offload experts under dynamic resource constraints with offline profiling and online scheduling. SiDA [20] dynamically load activated experts and offload inactivated experts according to hash tables. These proposals are mostly focused on MoE models. While helpful to reduce carbon footprint, we consider them to be orthogonal. Furthermore, naively implementing offloading without capacity planning would only cause more embodied and operational carbon due to high PCIE overhead.
Energy-aware Serving and Model Serving: A plethora of work in the space of model serving has emerged, focusing on providing SLOs with scheduling, placement, batching and autoscaling solutions. For example, AlpaServe [40], Clockwork [29], Sheperd [73] and Clipper [14] all are focusing on non-autoregreesive models. However, such general systems often overlook the distinct phases in LLM serving or autoregressive nature of LLMs. InFaaS [57] proposes a way to do model selection and autoscaling. More recently, userve [54] proposed power-aware LLM serving. More recently, Splitwise propose pd disaggregation for optimizing for cost, power, and throughput under SLOs  [53]. These general systems are not carbon-aware. Some newer works like Clover shows the tradeoffs between accuracy and carbon [39]. We show in our experiments that carbon-opt is different from energy-opt, and propose co-design solutions for both resource provision and runtime scheduling, leaning on the insights of asymmetric carbon cost of CPU and GPUs, without hurting model quality.
VIII Conclusion
Our work presents a framework for quantifying AI carbon footprints by integrating operational and embodied emissions, enabling AI sustainability optimization across various system stacks. It lays the groundwork for carbon-aware models, system-hardware co-design, and better hardware refresh strategies. We urge the research community to advance sustainable AI through holistic, quantitative approaches to environmental impact assessment and mitigation.